OpenStudy (anonymous):

Use the hang time function of V(T) = 48T2 to solve the following Tracy McGrady of the Houston Rockets, has a vertical leap of 40 inches. What is his hang Time? (10pts) Tip: V(T) is 40, so what does T equal?

OpenStudy (dumbcow):

solve for t 48t^2 = 40 t^2 = 40/48 = 5/6 square root both sides t = sqrt(5/6)